# Unleashing Multi-Hop Reasoning Potential in Large Language Models through Repetition of Misordered Context

### v1

@@ 4.3 Analysis
Does CoRe really identify the optimal order?
We conduct additional analysis to investigate whether the performance improvement from CoRe is indeed related to the order of the documents. Similar to the experiments described in Section 3.1, we permute the order of the documents in the context across all possible combinations and measure the model’s performance for each permuted context. Specifically, for 200 $k$-hop samples from MuSiQue, we construct a context with 5 documents, including the $k$ supporting documents and some random $5-k$ noisy documents assigned to each sample. As a result, each sample has 120 permuted contexts, and we sort the performance of each context from the worst order to the best order, based on $p_{\theta}(a\mid Q,C,I)$. Figure 5 shows the results for Llama. We plot the worst order with a red line and the best with a purple line. For other cases, we plot their performance with the line of a continuous gradient of colors from red to purple, corresponding to the performance range from the worst to the best case. In other words, the closer the color is to red in the rainbow spectrum, the lower $p_{\theta}(a\mid Q,C,I)$ is, and the closer to purple, the higher $p_{\theta}(a\mid Q,C,I)$ is. Across all query types, we find compelling evidence that CoRe contributes to performance improvements in terms of document order: The closer the context is to the worst order case, the greater the performance improvement from repetition. Additionally, we observe that as the number of hops increases, performance improves progressively with iterative repetition, indicating that multiple repetitions are more effective when the required reasoning steps increase. Appendix D shows the results for other LLMs in this analysis.
@@
We further analyze the impact of noisy documents on CoRe. In the synthetic task, we measure accuracy by increasing the number of repetitions in cases where the number of lists is 6, 3, and 1, respectively. Here, the decrease in the number of lists means a decrease in the proportion of information unrelated to the search target. Figure 6 shows the results of this analysis. Considering that the length of the context decreases as the number of lists handled in the context decreases, we perform more repetitions in cases with fewer lists. As a result, we confirm that as the proportion of noisy information within the context decreases, the performance degradation caused by excessive repetition is reduced. In particular, when there is no noisy information at all, performance did not decrease even with a large number of repetitions. Since noisy information adversely affects model reasoning at high repetition counts, exploring the valid number of repetitions required for model reasoning could be a potentially interesting future work based on our findings.
@@ 5.1 Robustness to Positional Bias
We analyzed the positional bias in the context based on the position of supporting documents. Here, positional bias refers to the phenomenon where the QA performance of an LLM varies depending on the position of the supporting documents within the context. This phenomenon is commonly known as the "lost-in-the-middle" problem Liu et al. (2023). We concatenated $k$ supporting documents in MuSiQue and measured the F1 scores of the baseline and CoRe while shifting the position of the documents from 0 to 18 by increments of 2. As shown in 7, we can observe that the performance gap between the baseline and CoRe increases as the position of the supporting documents moves toward the middle. Particularly for Llama (blue line) and Phi (green line), while the baseline suffers from the "lost-in-the-middle" problem, CoRe demonstrates robustness against positional bias, exhibiting minimal performance variation depending on the position of the supporting documents. Based on these results, we propose CoRe as a solution to mitigate the "lost-in-the-middle" problem in long context QA.
@@ 5.2 Enhancement in Retrieve-and-Reason
We evaluate the effectiveness of CoRe in a retrieve-and-reason task, which closely mirrors real-world scenarios. We utilize Contriever (Izacard et al., 2021) as the retriever model, constructing the context with up to 100 retrieved documents from MuSiQue samples. We use Llama for this experiment. We first assess CoRe’s performance with (1) the standard RAG, which directly retrieves the top 100 relevant documents to form the context. We also evaluate CoRe with two additional methods: (2) IRCoT (Trivedi et al., 2023), where retrieval is performed sequentially during the CoT reasoning process, and (3) Decompose, where the query is decomposed into sub-queries, and reasoning is conducted by deriving intermediate answers for each sub-query (Press et al., 2023). As shown in Table 2, CoRe significantly improves performance across all methods, demonstrating its effectiveness when integrated into various retrieval-based reasoning approaches. Appendix E presents the experimental details of this section.
@@
Table 2: F1 results for retrieve-and-reason task in MuSiQue. Dec. refers to Decompose
@@ 6 Conclusion
In this work, we introduce the misordered context problem, which has a critical impact on LLMs’ multi-hop reasoning capabilities. After formalizing this problem in terms of context, we propose a simple yet highly effective solution: the CoRe method. We demonstrate the contributions of our approach across various multi-hop reasoning benchmarks, including synthetic tasks, and elucidate its effectiveness through further studies. Our work highlights a key issue crucial for future discussions in multi-hop reasoning tasks with LLMs.
@@ Appendix E Experimental Details of Retrieve-and-Reason Task
In the question decomposition setting, before answering the query, the LLM is instructed to decompose the query into several sub-queries. Once the sub-queries are generated, the LLM is directed to perform step-by-step reasoning by following each sub-query, with the process guided by an instruction that ensures the sub-queries are considered during the answering phase. This setting aligns with the prompt design of Press et al. (2023). Table 8 shows the prompt template of query decomposition.
@@ Appendix F Ablation Studies
In this section, we conduct an ablation study to analyze the effectiveness of the CoRe methodology. Specifically, we investigate the impact of changing the position where CoRe is applied within the prompt. In our default setup, CoRe is applied through the assistant role, where the context is repeated in the responses generated by the assistant. However, we also consider an alternative setup where the context is repeated in the user role. We evaluate the performance of Llama, Mistral, Qwen, and Phi on 2WikiMultihopQA using F1 scores.
@@
Additionally, we conduct an ablation study to explore the effects of modifying the repeated context. Specifically, we examine 4 cases where the content of the context is altered during repetition: (1) each documents are paraphrased, (2) each documents are summarized, (3) the order of the documents is shuffled randomly, and (4) the order of the documents is reversed. We evaluate the performance of each of these cases. The experiments are conducted on 2WikiMultihopQA using Llama, where the paraphrasing and summarizing are directly performed by Llama responsible for reasoning.
@@ Appendix G Computational Environment
For the main experiments, we utilize NVIDIA A40 GPU. For the experiments for the retrieve-and-reason task, we utilize NVIDIA H100 80GB HBM3 GPU.

### v2

@@ 4.3 Analysis
Does CoRe really guide the model along the better reasoning chain?
We conduct additional analysis to investigate whether the performance improvement from CoRe is indeed related to the order of the documents. Similar to the experiments described in Section 3.1, we permute the order of the documents in the context across all possible combinations and measure the model’s performance for each permuted context. Specifically, for 200 $k$-hop samples from MuSiQue, we construct a context with five documents, including the $k$ supporting documents and some random $5-k$ noisy documents assigned to each sample. As a result, each sample has 120 permuted contexts, and we sort the performance of each context from the worst order to the best order, based on $p_{\theta}(a\mid Q,C,I)$. Figure 5 shows the results for Llama. We plot the worst order with a red line and the best with a purple line. For other cases, we plot their performance with the line of a continuous gradient of colors from red to purple, corresponding to the performance range from the worst to the best case. In other words, the closer the color is to red in the rainbow spectrum, the lower $p_{\theta}(a\mid Q,C,I)$ is, and the closer to purple, the higher $p_{\theta}(a\mid Q,C,I)$ is. Across all query types, we find compelling evidence that CoRe contributes to performance improvements in terms of document order: The closer the context is to the worst order case, the greater the performance improvement from repetition. Additionally, we observe that as the number of hops increases, performance improves progressively with iterative repetition, indicating that multiple repetitions are more effective when the required reasoning steps increase. Appendix D shows the results for other LLMs in this analysis.
@@
We further analyze the impact of noisy documents on CoRe. In the synthetic task, we measure accuracy by increasing the number of repetitions in cases where the number of lists is 6, 3, and 1, respectively. Here, the decrease in the number of lists means a reduction in the proportion of information unrelated to the search target. Figure 6 shows the results of this analysis. Considering that the length of the context decreases as the number of lists handled in the context decreases, we perform more repetitions in cases with fewer lists. As a result, we confirm that as the proportion of noisy information within the context decreases, the performance degradation caused by excessive repetition is reduced. In particular, when there is no noisy information, performance did not decrease even with a large number of repetitions. Since noisy information adversely affects model reasoning at high repetition counts, exploring the valid number of repetitions required for model reasoning could be an interesting future work based on our findings.
@@ 5.1 Robustness to Positional Bias
We analyze the absolute positional bias in the context based on the position of supporting documents. Here, positional bias refers to the phenomenon where the QA performance of an LLM varies depending on the position of the supporting documents within the context, known as the "lost-in-the-middle" problem Liu et al. (2023). We concatenate $k$ supporting documents in MuSiQue and measure the F1 scores of the baseline without repetition ($\hat{k}=1$) and CoRe ($\hat{k}=2$) while shifting the position of the documents from 0 to 18 by increments of 2. As shown in 7, we observe that the performance gap between the baseline and CoRe increases as the position of the supporting documents moves toward the middle. Particularly for Llama (blue line) and Phi (green line), while the baseline suffers from the "lost-in-the-middle" problem, CoRe demonstrates robustness against positional bias, exhibiting minimal performance variation depending on the position of the supporting documents. Based on these results, we demonstrate that CoRe is a solution to mitigate the "lost-in-the-middle" problem.
@@ 5.2 Enhancement in Retrieve-and-Reason
We evaluate the effectiveness of CoRe ($\hat{k}=2$) in a retrieve-and-reason task, which closely mirrors real-world scenarios. We utilize Contriever (Izacard et al., 2021) as the retriever model, constructing the context with up to 100 retrieved documents from MuSiQue samples. We use Llama for this experiment. We first assess CoRe’s performance with (1) the standard RAG, which directly retrieves the top 100 relevant documents to form the context. We also evaluate CoRe with two additional methods: (2) IRCoT (Trivedi et al., 2023), where retrieval is performed sequentially during the CoT reasoning process, and (3) Decompose, where the query is decomposed into sub-queries, and reasoning is conducted by deriving intermediate answers for each sub-query (Press et al., 2023). As shown in Table 2, CoRe significantly improves performance across all methods, demonstrating its effectiveness when integrated into various retrieval-based reasoning approaches. Appendix E presents the experimental details of this section.
@@
Table 2: F1 results for retrieve-and-reason tasks in MuSiQue. Dec. refers to Decompose method.
@@ 6 Conclusion
In this work, we introduce the misordered context problem, which critically impacts LLMs’ multi-hop reasoning capabilities. Formalizing this problem regarding context augmentation, we propose a simple yet highly effective solution: the CoRe method, which optimally arranges contiguous reasoning segments within the context to enhance performance. We demonstrate the contributions of our approach across various multi-hop reasoning benchmarks, including synthetic tasks, and elucidate its effectiveness through further studies. Our work highlights a key issue crucial for future discussions in multi-hop reasoning tasks with LLMs. Furthermore, we believe CoRe could provide meaningful insights for research on the test-time scaling of LLMs that perform complex reasoning.
@@
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University, Samsung Electronics Co., Ltd (IO240311-09242-01), and a grant from the Yang Young Foundation.
@@ Appendix E Experimental Details of Retrieve-and-Reason Task
In the question decomposition setting, LLM is instructed to decompose the query into several sub-queries before answering the query. Once the sub-queries are generated, the LLM is directed to perform step-by-step reasoning by following each sub-query, with the process guided by an instruction that ensures the sub-queries are considered during the answering phase. This setting aligns with the prompt design of Press et al. (2023). Table 8 shows the prompt template of query decomposition.
@@ Appendix F Ablation Studies
F.1 Prompt Style Ablation
In this section, we investigate the impact of changing the position where CoRe is applied within the prompt. Specifically, we conduct an ablation study for the role where CoRe is applied in the chat template prompt. In our default setup, CoRe is applied through the assistant role, where the context is repeated in the responses generated by the assistant. However, we also consider an alternative setup where the context is repeated in the user role. We evaluate the performance of Llama, Mistral, Qwen, and Phi on 2WikiMultihopQA using F1 scores. $\hat{k}=2$ for CoRe in this study.
@@ F.1 Prompt Style Ablation
F.2 Repetition Style Ablation
Additionally, we conduct an ablation study to explore the effects of modifying the repeated context. Specifically, we examine 4 cases where the content of the context is altered during repetition: (1) each documents are paraphrased, (2) each documents are summarized, (3) the order of the documents is shuffled randomly, and (4) the order of the documents is reversed. We evaluate the performance of each of these cases. The experiments are conducted on 2WikiMultihopQA using Llama, where the paraphrasing and summarizing are directly performed by Llama responsible for reasoning. $\hat{k}=2$ for CoRe in this study.
@@ Appendix G Computational Environment
For the main experiments, we utilize NVIDIA A40 GPU. For the experiments for the retrieve-and-reason task, we utilize NVIDIA H100 80GB HBM3 GPU. All the models except GPT-4o-mini in our work are from huggingface. We utilize GPT-4o-mini through OpenAI API package.